# BiDoRA: Bi-level Optimization-Based Weight-Decomposed Low-Rank Adaptation

### v1

@@ Abstract
Parameter-efficient fine-tuning (PEFT) of large language models (LLMs) has gained considerable attention as a flexible and efficient way of adapting LLMs to downstream tasks. Among these methods, weighted decomposed low-rank adaptation (DoRA) has emerged as a promising approach. DoRA bridges the gap between low-rank adaptation (LoRA) and full fine-tuning (FT) by decomposing the weight matrices into magnitude and direction components, thereby maintaining learning behavior similar to FT. Although DoRA shows encouraging performance, it introduces additional parameters compared to LoRA, which potentially increases the risk of overfitting. Moreover, optimizing magnitude and direction simultaneously leads to a coupled gradient updating pattern for both components, limiting its learning capacity. To overcome these limitations, we propose BiDoRA, a bi-level optimization-based PEFT method. In BiDoRA, the direction and magnitude components are optimized on two distinct datasets at different optimization levels, mitigating the risk of overfitting. Additionally, the asynchronous optimization of the two components promotes their decoupling, allowing for more flexible gradient updates suitable for various downstream tasks. Evaluation of BiDoRA on fourteen datasets spanning natural language understanding, natural language generation, and token classification reveals that it significantly outperforms DoRA and other PEFT methods. The superior performance of BiDoRA underscores its effectiveness. The code for BiDoRA is available at https://anonymous.4open.science/r/BiDoRA-5D31.
@@ 1 Introduction
Large language models (LLMs) (Radford et al., 2019; Brown et al., 2020) have achieved state-of-the-art results across a broad range of NLP tasks, from natural language understanding (NLU) (Wang et al., 2019) to natural language generation (NLG) (Novikova et al., 2017). Although LLMs are pre-trained on extensive datasets in the general domain, additional methods are required to adapt them to specific downstream datasets for better performance. For example, full fine-tuning (FT) involves updating all pre-trained parameters on downstream datasets (Qiu et al., 2020), which typically yields superior results. However, as LLMs continue to scale, FT incurs significant computational costs and increases the risk of overfitting during the fine-tuning process (Karimi Mahabadi et al., 2021). Alternatively, in-context learning adapts LLMs to new tasks by incorporating a few data examples directly into the prompt, eliminating the need for parameter updates (Brown et al., 2020). While this approach substantially reduces computational costs and mitigates the risk of overfitting, it is constrained by the use of only a limited number of examples from the downstream data, often leading to suboptimal performance.
To address the limitations above, parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Hu et al., 2021) have been introduced as a promising solution. PEFT approaches update only a subset of the pre-trained parameters, achieving performance comparable to FT while requiring significantly fewer computational resources. Among these methods, low-rank adaptation (LoRA, Hu et al. (2021)) has gained popularity due to its simplicity and effectiveness. LoRA attaches low-rank matrices to the pre-trained model weights, updating only these matrices during fine-tuning without changing the model architecture. However, LoRA exhibits a parameter updating pattern distinct from FT, which may limit its learning capacity (Liu et al., 2024). Specifically, when decomposing the parameter updates of both methods into magnitude and direction components, the correlation between these components tends to be positive in LoRA, whereas it is negative in FT. To bridge this gap, weight-decomposed low-rank adaptation (DoRA, Liu et al. (2024)) introduces an explicit reparameterization of the pre-trained weights matrix as the product of magnitude and direction components. The model can be trained end-to-end in the same way, with direction components parameterized with incremental low-rank matrices. This approach enables DoRA to share similar learning patterns with FT, thereby outperforming LoRA in multiple tasks. Nonetheless, DoRA introduces additional parameters compared to LoRA, which can exacerbate overfitting issues, particularly when adapting to smaller downstream datasets (See Figure 2). Furthermore, in DoRA, the magnitude and incremental direction components are optimized concurrently, leading to a highly constrained updating pattern that may overlook the diverse learning patterns required for different downstream tasks.
@@
Figure 1: An overview of BiDoRA. BiDoRA performs PEFT using a bi-level optimization framework. At the lower level, BiDoRA learns the incremental direction component $\Delta V$ of the update matrices using the training split of the downstream dataset. At the upper level, BiDoRA optimizes the magnitude component $m$ with optimized $\Delta V$ from the lower level, using the validation split of the dataset. After determining the optimal magnitude, the direction component undergoes further fine-tuning on a combined set of both training and validation splits to maximize overall performance.
@@
To address the challenges above, we propose BiDoRA, a bi-level optimization-based weight-decomposed low-rank adaptation method for parameter-efficient fine-tuning of LLMs. BiDoRA mitigates overfitting and facilitates a flexible weight updating pattern by separately optimizing the magnitude and incremental direction components on different splits of the downstream datasets using distinct optimization steps. BiDoRA is based on a bi-level optimization framework: At the lower level, the low-rank incremental direction component is updated using the training split of the downstream dataset, while the magnitude component remains fixed. At the upper level, the magnitude component is updated by minimizing the loss on the validation split via hypergradient descent. These two optimization steps are performed iteratively until convergence. Figure 1 provides an overview of BiDoRA.
@@
In BiDoRA, the two distinct components are trained on separate splits of the dataset, which effectively reduces the risk of overfitting. A similar strategy is utilized in the well-established practice of differentiable neural architecture search (DARTS, Liu et al. (2018)), where architecture and model parameters are learned using different splits of the training dataset. Since the architecture selection module introduces additional parameters beyond the base model, directly optimizing this overly complex supernetwork can result in severe overfitting. Similarly, by treating the magnitude component as the architecture and the incremental direction component as the model in neural architecture search, training these components on separate datasets helps reduce overfitting. As shown in Figure 2, BiDoRA demonstrates better resistance to overfitting compared to DoRA, given the smaller performance gap between the training set and test set. Furthermore, the asynchronous gradient update steps at the two optimization levels in BiDoRA facilitate better decoupling of the magnitude component from the incremental direction component, leading to a more flexible update pattern that closely resembles FT. As illustrated in Figure 3, the updates of the magnitude and direction components across different layers using BiDoRA have a correlation value that is closest to that of FT, highlighting its superior learning capability compared to both DoRA and LoRA.
@@
  • We propose BiDoRA, a novel parameter-efficient fine-tuning method based on bi-level optimization. In contrast to DoRA, which trains the magnitude and incremental direction components on a single dataset, BiDoRA optimizes these components on different splits of a downstream dataset through distinct optimization steps.
@@
  • Extensive experiments on fourteen datasets highlight the superior performance of BiDoRA. BiDoRA consistently surpasses several baseline methods, including LoRA and DoRA, across tasks such as text classification, language generation, and token classification.
Figure 2: Training and test accuracy versus global training steps on the ModHayes split of the Reuters21578 dataset when fine-tuning a RoBERTa-base model using DoRA and BiDoRA. The training and test curves for DoRA show a larger gap compared to BiDoRA, highlighting the effectiveness of our method in reducing overfitting.
(a) Full FT($k=-65.816$)
(b) LoRA($k=0.836$)
(c) DoRA($k=-1.784$)
(d) BiDoRA($k=-8.042$)
Figure 3: Magnitude and direction updates for (a) FT, (b) LoRA, (c) DoRA, and (d) BiDoRA of the query matrices across different layers and intermediate steps after fine-tuning the GPT2 model on the E2E dataset, where $k$ denotes the correlation value. Different markers represent matrices from different training steps, with each color corresponding to a specific layer.
@@ 2.1 Parameter Efficient fine-tuning Methods
Parameter-efficient fine-tuning (PEFT) methods aim to reduce the high costs associated with fully fine-tuning large-scale models by updating only a relatively small subset of pre-trained parameters, rather than the entire model, to adapt to downstream tasks. Existing PEFT methods can be mainly categorized into three types. The first category, known as adapter-based methods, injects additional trainable modules into the original frozen backbone. For instance, Houlsby et al. (2019) suggests adding linear modules in sequence to existing layers, while He et al. (2021) proposes integrating these modules in parallel with the original layers to enhance performance. The second category is prompt tuning methods, which add extra soft tokens (prompts) to the initial input. During the fine-tuning stage, only these trainable soft tokens are updated, as demonstrated in works such as Lester et al. (2021) and Razdaibiedina et al. (2023). Unfortunately, the first two categories lead to increased inference latency compared to fully fine-tuned models. The third category is low-rank adaptation methods, pioneered by the foundational work LoRA (Hu et al., 2021). These methods attach low-rank matrices to pre-trained weights and use only these matrices for weight updates during fine-tuning. Since low-rank updates can be merged with pre-trained weights before inference, low-rank adaptation-based PEFT methods do not increase inference time. Following LoRA, Zhang et al. (2023) applies SVD decomposition to low-rank matrices and prunes less significant singular values for more efficient updates. Zhang et al. (2024) uses meta-learning to search for the optimal rank of LoRA matrices, further improving its performance on downstream tasks. Most recently,Liu et al. (2024) uses weight decomposition analysis to reveal that LoRA exhibits a distinct weight updating pattern compared to FT, which may constrain its learning capacity. Therefore, DoRA (Liu et al., 2024) was then proposed to bridge the gap between LoRA and FT. DoRA decomposes the pre-trained weights into two components—magnitude and direction—and fine-tunes both, which results in a more closely aligned updating pattern compared to FT.
@@ 2.2 Bi-level Optimization
Bi-level optimization (BLO) has been widely applied in various machine learning tasks, including meta-learning (Finn et al., 2017; Rajeswaran et al., 2019), neural architecture search (Liu et al., 2018; Zhang et al., 2021), and hyperparameter optimization (Lorraine et al., 2020; Franceschi et al., 2017). Despite its wide usage, solving BLO problems can be challenging due to the inherent nature of nested optimization problems. Several algorithms have been proposed to address this challenge, including zeroth-order methods such as Bayesian optimization (Cui & Bai, 2019) and first-order algorithms based on hypergradients (Pearlmutter & Siskind, 2008; Lorraine et al., 2020). Among these approaches, gradient-based BLO has received significant attention because it can scale to high-dimensional problems with a large number of trainable parameters. In this work, we extend the application scenarios of gradient-based BLO to develop a robust and effective parameter-efficient fine-tuning method for pre-trained models.
3 Preliminary: Weight-Decomposed Low-Rank Adaptation
@@
LoRA (Hu et al., 2021) involves attaching the product of two low-rank matrices to the pre-trained weights and fine-tuning these low-rank matrices on downstream datasets with the pre-trained weights frozen. It is based on the assumption that parameter updates made during fine-tuning exhibit a low intrinsic rank. Formally, given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA attaches a low-rank update matrix $\Delta W\in\mathbb{R}^{d\times k}$ to the pre-trained weight. This update matrix can be further decomposed as $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ are two low-rank matrices, with $r\ll\min(d,k)$. Consequently, the weight matrix $W^{\prime}$ after LoRA fine-tuning can be represented as follows:
$W^{\prime}=W_{0}+\Delta W=W_{0}+BA$ (1)
@@
In this setup, the pre-trained weight matrix $W_{0}$ remains fixed during the fine-tuning process, while the LoRA matrix $\Delta W$ is updated. However, after performing weight decomposition on fine-tuned weight matrices, it was found that LoRA and full fine-tuning exhibit different learning patterns (Liu et al., 2024). To bridge this discrepancy, weight-decomposed low-rank adaptation (DoRA, Liu et al. (2024)) further reparameterizes the weight matrices by explicitly decomposing them into learnable magnitude and direction components. Formally, DoRA performs adaption as follows:
$W^{\prime}=m\frac{V+\Delta V}{\|V+\Delta V\|_{c}}=m\frac{W_{0}+BA}{\|W_{0}+BA\|_{c}}$ (2)
@@
where the incremental direction component $\Delta V$ is parameterized as a product of two learnable low-rank matrices, $B$ and $A$, while the magnitude component $m\in\mathbb{R}^{1\times k}$ is a learnable vector. Here, $\|\cdot\|_{c}$ represents the vector-wise norm of a matrix computed across each column. In DoRA, both components are optimized concurrently on a single downstream dataset.

### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) is a flexible and efficient method for adapting large language models (LLMs) to downstream tasks. Among these methods, weight-decomposed low-rank adaptation (DoRA) is a promising approach that decomposes weight matrices into magnitude and direction components to mimic full fine-tuning (FT) better. However, DoRA’s simultaneous optimization of these components makes it over-expressive, increases the risk of overfitting, and creates a coupled updating pattern that limits its learning capacity. To address these issues, we propose Bi-level Optimization-Based Weight-Decomposed Low-Rank Adaptation (BiDoRA), a novel PEFT method based on a bi-level optimization framework. BiDoRA fundamentally differs from DoRA by optimizing the magnitude and direction in two separate, asynchronous loops using distinct training and validation data splits. This decoupled optimization process effectively mitigates overfitting and allows for more flexible updates that align even more closely with FT. For instance, weight decomposition analysis shows BiDoRA achieves a magnitude-direction update correlation of $\mathbf{-8.042}$, significantly closer to the FT ideal compared to $\mathbf{-1.784}$ for DoRA. Evaluation of BiDoRA on diverse tasks spanning natural language understanding, generation, token classification, and extremely small biomedical datasets reveals that it consistently outperforms DoRA and a wide range of leading PEFT methods. This improvement is statistically significant, as demonstrated on the GLUE benchmark where BiDoRA surpasses DoRA with a p-value of $\mathbf{2.4\times 10^{-4}}$ in terms of the Wilcoxon signed-rank test. The code for BiDoRA is available at https://github.com/t2ance/BiDoRA.
@@ 1 Introduction
Figure 1: An overview of BiDoRA. BiDoRA performs PEFT using a BLO framework. At the lower level, BiDoRA learns the direction component $\Delta\mathbf{V}$ of the update matrices using the training split of the downstream dataset. At the upper level, BiDoRA optimizes the magnitude component $\mathbf{m}$ with optimized $\Delta\mathbf{V}$ from the lower level, using the validation split of the dataset. After determining the optimal magnitude, the direction component undergoes further fine-tuning on a combined set of both training and validation splits to maximize overall performance.
Large language models (LLMs) (Radford et al., 2019; Brown et al., 2020) have achieved state-of-the-art results across a broad range of NLP tasks, from natural language understanding (NLU) (Wang et al., 2019) to natural language generation (NLG) (Novikova et al., 2017). Parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Hu et al., 2022b) have been introduced as a promising solution for adapting LLMs for downstream data. PEFT approaches update only a subset of the pre-trained parameters, achieving performance comparable to full-finetuning (FT) while requiring significantly fewer computational resources.
One popular type of PEFT is low-rank adaptation (LoRA, Hu et al. (2022b)), which attaches low-rank matrices to the pre-trained weights and updates only these matrices during fine-tuning. Liu et al. (2024a) shows that when decomposing the weights into magnitude and direction, their correlation (See Appendix˜D) tends to be positive in LoRA, whereas it is negative in FT. To bridge the training pattern distinction, they introduce an explicit reparameterization of the pre-trained weights matrix. The method, named DoRA, decomposes the weights into the column-wise product of magnitude and direction, which determines the direction and magnitude of the weight update, respectively. This approach enables DoRA to share similar learning patterns with FT, thereby outperforming LoRA in multiple tasks. Nonetheless, DoRA introduces additional parameters and over-expressive architecture compared to LoRA, which can exacerbate overfitting issues when adapting to small downstream datasets (See Table˜3). Furthermore, in DoRA, the magnitude and direction components are optimized concurrently, leading to a constrained updating pattern due to shared optimization setup (e.g., learning rate, optimizer, batch size).
@@
To address the challenges above, we propose BiDoRA, a Bi-level Optimization-Based Weight-Decomposed Low-Rank Adaptation method for PEFT. BiDoRA facilitates an even more flexible updating pattern and mitigates overfitting by separately optimizing the two components on different data splits with distinct optimization levels. BiDoRA is based on a bi-level optimization (BLO) framework: At the lower level, the low-rank direction component is updated using the training split, while the magnitude component remains fixed. At the upper level, the magnitude component is updated by minimizing the loss on the validation split via hypergradient descent. Subsequently, the direction component is further fine-tuned with the optimal magnitude frozen to maximize the performance. These two optimization steps are performed iteratively until convergence. Fig.˜1 provides an overview of BiDoRA.
@@
A similar strategy of combating overfitting based on BLO has been utilized in the well-established practice of differentiable neural architecture search (DARTS, Liu et al. (2019)), where architecture and sub-networks are learned using different dataset splits. Optimizing the selection variables and sub-networks in a single loop can result in an over-expressive network since the selection variables tend to select all sub-networks to achieve the best expressiveness, which, however, incurs severe overfitting. In contrast, training the sub-networks with the selection module fixed on the training split while validating the effectiveness of the selection module on the unseen validation split effectively eliminates the risk of overfitting. Similarly, we treat the magnitude component as the architecture and the direction component as the sub-networks and train these components on separate datasets. As shown in Table˜3, BiDoRA demonstrates better resistance to overfitting compared to DoRA, given the smaller performance gap between the training set and test set. Furthermore, the asynchronous gradient update steps at the two optimization levels in BiDoRA facilitate better decoupling of the two components, leading to a more flexible update pattern that closely resembles FT. As illustrated in Fig.˜3, the updates across different layers using BiDoRA have a correlation value that is closest to that of FT, highlighting its superior learning capability compared to both DoRA and LoRA.
@@
  • We propose BiDoRA, a novel PEFT method based on bi-level optimization. In contrast to DoRA, which trains the magnitude and direction components on a single dataset, BiDoRA optimizes these components at different optimization levels.
@@
  • Extensive experiments on various downstream tasks highlight the superior performance of BiDoRA. BiDoRA consistently surpasses several baseline methods, including LoRA and DoRA.
@@ 2.1 Parameter Efficient Fine-Tuning Methods
Parameter-efficient fine-tuning (PEFT) methods aim to reduce the high costs associated with full fine-tuning large-scale models by updating only a relatively small subset of pre-trained parameters, rather than the entire model, to adapt to downstream tasks. Existing PEFT methods can be mainly categorized into three types.
The first category, known as adapter-based methods, injects additional trainable modules into the original frozen backbone. For instance, Houlsby et al. (2019) suggests adding linear modules in sequence to existing layers, while He et al. (2022) proposes integrating these modules in parallel with the original layers to enhance performance. Recent advances include SAN (Xu et al., 2023), FADA (Bi et al., 2024), and SET (Yi et al., 2024). SAN presents a side adapter network attached to a frozen CLIP model, which contains two branches for predicting mask proposals and attention biases. FADA introduces a frequency-adapted learning scheme that uses the Haar wavelet transform to decompose frozen features into low- and high-frequency components, which are processed separately to enhance domain generalization. SET proposes a spectral-decomposed token learning framework that leverages the Fast Fourier Transform to separate frozen features into amplitude and phase components, enhancing them with spectral tokens and attention optimization.
The second category is prompt tuning methods, which add extra soft tokens (prompts) to the initial input. During the fine-tuning stage, only these trainable soft tokens are updated, as demonstrated in works such as Lester et al. (2021) and Razdaibiedina et al. (2023). Unfortunately, the first two categories lead to increased inference latency compared to fully fine-tuned models.
The third prominent category focuses on low-rank adaptation, pioneered by LoRA (Hu et al., 2022a). LoRA injects trainable, low-rank matrices into a model’s layers, freezing the original weights. A key advantage is that these low-rank updates can be merged into the original weights before inference, thus incurring no additional latency. Subsequent works have aimed to improve LoRA’s efficiency and performance. For instance, AdaLoRA (Zhang et al., 2023) dynamically reallocates the parameter budget based on the importance scores of weight matrices. Zhang et al. (2024b) uses meta-learning to search for the optimal rank of LoRA matrices, further improving its performance on downstream tasks. Pushing parameter efficiency further, VeRA (Kopiczko et al., 2024) employs a single pair of shared low-rank matrices across all layers, while AFLoRA (Liu et al., 2024b) freezes a portion of adaptation parameters based on a learned score. A distinct sub-direction has emerged that performs adaptation in the frequency domain, including FourierFT (Gao et al., 2024), LaMDA (Azizi et al., 2024), SSH (Shen et al., 2025b), and MaCP (Shen et al., 2025a). These methods learn updates in transformed spectral spaces, such as the Fourier, discrete Hartley, or discrete cosine domains, rather than directly in the weight space. Other research has focused on bridging the performance gap between LoRA and full fine-tuning. Liu et al. (2024a) found that LoRA’s update patterns differ significantly from full fine-tuning, potentially constraining its learning capacity. To mitigate this, they proposed DoRA (Liu et al., 2024a), which decomposes pre-trained weights into magnitude and direction components and uses LoRA for efficient directional updates, better mimicking full fine-tuning.
@@ 2.2 Bi-level Optimization
Bi-level optimization (BLO) has been widely applied in various machine learning tasks, including meta-learning (Finn et al., 2017; Rajeswaran et al., 2019), neural architecture search (NAS) (Liu et al., 2019; Zhang et al., 2021), and hyperparameter optimization (Lorraine et al., 2020; Franceschi et al., 2017). Despite its wide usage, solving BLO problems can be challenging due to the inherent nature of nested optimization problems. Several algorithms have been proposed to address this challenge, including zeroth-order methods such as Bayesian optimization (Cui & Bai, 2019) and first-order algorithms based on hypergradients (Pearlmutter & Siskind, 2008; Lorraine et al., 2020). Among these approaches, gradient-based BLO has received significant attention because it can scale to high-dimensional problems with a large number of trainable parameters.
Inspired by NAS, where a bi-level approach is used to learn an architecture and its sub-network weights on separate data splits to prevent overfitting, we adapt the BLO framework to parameter-efficient fine-tuning (PEFT), specifically for the weight-decomposed adaptation introduced by DoRA. Unlike in NAS, where BLO searches for a network architecture, BiDoRA repurposes it to decouple the optimization of a weight matrix’s two components: magnitude and direction. This approach marks a significant departure from previous PEFT methods like LoRA and DoRA, which optimize all trainable parameters simultaneously on a single dataset. In this work, we extend the application of gradient-based BLO to develop a robust and effective PEFT method for pre-trained models. By assigning the magnitude and direction components to different optimization levels with distinct data splits, BiDoRA creates a decoupled, flexible updating pattern that better mitigates overfitting and more closely resembles the learning behavior of full fine-tuning.
3 Preliminary
@@
LoRA (Hu et al., 2022b) involves attaching the product of two low-rank matrices to the pre-trained weights and fine-tuning these low-rank matrices on downstream datasets with the pre-trained weights frozen. It is based on the assumption that parameter updates made during fine-tuning exhibit a low intrinsic rank. Formally, given a pre-trained weight matrix $\mathbf{W_{0}}\in\mathbb{R}^{d\times k}$, LoRA attaches a low-rank update matrix $\Delta\mathbf{W}\in\mathbb{R}^{d\times k}$ to the pre-trained weight. This update matrix can be decomposed as $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times k}$ are two low-rank matrices, with $r\ll\min(d,k)$. Consequently, the weight matrix $\mathbf{W}^{\prime}$ is represented as follows:
$\mathbf{W}^{\prime}=\mathbf{W_{0}}+\Delta\mathbf{W}=\mathbf{W_{0}}+\mathbf{B}\mathbf{A}$ (1)
@@
In this setup, only the LoRA matrix $\Delta\mathbf{W}$ is updated. Liu et al. (2024a) found that LoRA and full fine-tuning exhibit different learning patterns by performing weight decomposition on fine-tuned weight matrices (See Appendix˜D). To bridge this discrepancy, weight-decomposed low-rank adaptation (DoRA, Liu et al. (2024a)) further reparameterizes the weight matrices by explicitly decomposing them into learnable magnitude and direction components. Formally, DoRA performs adaptation as follows:
$\mathbf{W}^{\prime}=\mathbf{m}\frac{\mathbf{V}+\Delta\mathbf{V}}{\|\mathbf{V}+\Delta\mathbf{V}\|_{c}}=\mathbf{m}\frac{\mathbf{W_{0}}+\mathbf{B}\mathbf{A}}{\|\mathbf{W_{0}}+\mathbf{B}\mathbf{A}\|_{c}}$ (2)
@@
where $\Delta\mathbf{V}$ is a product of two learnable low-rank matrices, $\mathbf{B}$ and $\mathbf{A}$, while the magnitude component $\mathbf{m}\in\mathbb{R}^{1\times k}$ is a learnable vector. Here, $\|\cdot\|_{c}$ represents the vector-wise norm of a matrix computed across each column, using the $L_{2}$ norm. In DoRA, both components are optimized concurrently on a single downstream dataset. In this work, we aim to improve DoRA by further decoupling the training of the two components.